# LayeredMAPF: a decomposition of MAPF instance without compromising solvability

### v1

@@ II-D2 Lazy constraints addition for MAPF (LaCAM)
It is noteworthy that when LaCAM encounters a known configuration, it reinserts the corresponding high-level node into the open set (which stores nodes for high-level search). This action prevents repeated appearance of configurations, thereby preventing LaCAM from accepting external paths as dynamic obstacles to avoid, as all agents may need to wait at times to avoid conflicts with external paths.
@@
Okumura et al. proposed LaCAM2 [okumura2023improving], which introduces two enhancements to LaCAM. The first enhancement is its anytime version, called LaCAM*, which converges to optimality eventually, provided that solution costs accumulate transition costs. The second enhancement adds a swap action to the successor generator of LaCAM, enabling the quick generation of initial solutions.
III Preliminaries
III-A Basic definitions
@@
In this section, we present fundamental definitions of multi-agent pathfinding. Since our method is dimension-independent, the provided definitions apply to any-dimensional grid space $\mathcal{C}_{\mathcal{N}}$.
III-A1 Grid space
@@
Let $\mathcal{C}_{\mathcal{N}}$ ($\mathcal{N}\geq 2$) denote a finite $\mathcal{N}$-dimensional integer Euclidean space, where the size of the space is defined as $\mathcal{D}\rightarrow\mathcal{C}_{\mathcal{N}}$, with $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{i},\ldots,d_{\mathcal{N}}\}$ and $d_{i}\in\mathbb{N}$. The coordinates of an element g in this space are defined as a vector $(x_{1},x_{2},\ldots,x_{i},\ldots,x_{\mathcal{N}})$, where $x_{i}\in([0,d_{i})\cap\mathbb{N})$.
III-A2 Grid states
@@
There are only two possible states for a grid/element in $\mathcal{C}_{\mathcal{N}}$: passable or unpassable. The set of all passable grids in $\mathcal{C}_{\mathcal{N}}$ is denoted as $\mathcal{F}\rightarrow\mathcal{C}_{\mathcal{N}}$, while the set of all unpassable grids is denoted as $\mathcal{O}\rightarrow\mathcal{C}_{\mathcal{N}}$. Therefore, $(\mathcal{F}\rightarrow\mathcal{C}_{\mathcal{N}})\cup(\mathcal{O}\rightarrow\mathcal{C}_{\mathcal{N}})=\mathcal{C}_{\mathcal{N}}$, $(\mathcal{F}\rightarrow\mathcal{C}_{\mathcal{N}})\cap(\mathcal{O}\rightarrow\mathcal{C}_{\mathcal{N}})=\emptyset$.
For convenience, we denote all grids outside of $\mathcal{C}_{\mathcal{N}}$ as unpassable. If a path passes no unpassable grids, we denote it as collision-free; otherwise, it collides with obstacles.
III-A3 Neighborhood of grid
Denote $\delta(g)$ as the neighborhood of grid $g$, which consists of the closest $2\mathcal{N}$ grids near $g$.
III-A4 Agents
@@
Assuming there are $k$ agents $A=\{a_{1},a_{2},\ldots,a_{k}\}$ in the grid space $\mathcal{C}_{\mathcal{N}}$, where each agent always occupies a passable grid. Their starting grids are denoted by $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{k}\}$, and their target grids are denoted by $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{k}\}$.
@@
Each agent has a unique starting grid and a targeting grid. Specifically, for all $i\in\{1,2,\ldots,k\}$ and $j\in\{1,2,\ldots,k\}$ with $i\neq j$, we have $s_{i}\neq s_{j}$, $t_{i}\neq t_{j}$, $s_{i}\neq t_{j}$, and $t_{i}\neq s_{j}$.
III-A5 Time stamp
@@
Time is discretized into timesteps. At each timestep, every agent can either move to an adjacent grid or wait at its current grid.
III-A6 Path
@@
A path $p_{i}$ for agent $a_{i}$ is a sequence of grids that are pairwise adjacent or identical (indicating a wait action), starting at the start grid $s_{i}$ and ending at the target grid $t_{i}$. The cost of a path is the number of time steps it takes, i.e., the size of the path sequence $|p_{i}|$. We assume that the agents remain at their targets forever after completing their paths.
III-A7 Conflict
@@
We say that two agents have a conflict if they are at the same grid or exchange their positions at the same timestep.
III-A8 Solution
@@ III-A9 Solvability
A method to determine whether a MAPF instance is solvable is whether every agent can find a collision-free path (i.e., pass no unpassable grids) from the grid space. It is noteworthy that passing through other agents’ starting or target grids is acceptable.
III-A10 Decomposition of MAPF instance
Considering that the cost of solving a MAPF instance grows nearly exponentially as the number of agents increases [li2022mapf], decomposing a MAPF instance and solving subproblems typically incurs lower costs compared to solving the raw instance.
@@
To reduce the cost of solve MAPF instance, we proposed a new technique: splitting a MAPF instance with multiple agents into multiple subproblems, each with fewer agents but under the same map, and solving these subproblems in a specific order. Here, we define the number of agents in each subproblem as the size of the subproblem. When solving a subproblem, it is crucial to ensure that the current paths do not conflict with the paths from previous subproblems. If every subproblem is solvable, combining all subproblems’ results yields the solution of the original MAPF instance.
@@
The criteria for judging whether one decomposition is better than another involves sorting the subproblems’ sizes in decreasing order and comparing the sizes of the subproblems from the largest to the smallest. The first smaller subproblem encountered indicates the better decomposition.
@@
For example, suppose there are two decompositions of a MAPF instance with 100 agents. The sizes of the subproblems for the first decomposition are 40, 20, 15, 14, and 11, while for the second decomposition, they are 40, 20, 19, 13, and 8. The first decomposition is considered better than the second one because the third cluster’s size (15) in the first decomposition is smaller than the second decomposition’s third subproblem size (19).
@@
We compare the quality of decomposition in this manner because the cost of solving a MAPF instance grows exponentially as the number of agents increases [li2022mapf]. In theory, solving all agents together has a higher chance of finding a shorter solution compared to solving them separately. Therefore, MAPF methods with decomposition of MAPF instances tend to produce worse solutions than raw methods. However, we provide no theoretical analysis regarding the extent of this degradation in this manuscript; hence, we conduct empirical comparisons through extensive testing in Section VI.
In implementation, we measure the quality of decomposition by calculating the ratio between the size of the maximum subproblem and the number of agents of the raw MAPF instance. Since the cost of solving a MAPF instance is mainly determined by the maximum subproblem’s size, we refer to such a ratio as the decomposition rate in the following text.
Based on the criteria for comparing decompositions, a MAPF instance has a certain optimal decomposition that is better than all other decompositions. We have not found a direct way to obtain the optimal decomposition, but we can optimize an initial decomposition progressively to achieve a better decomposition.
III-B Theory basis of decomposition
A decomposition of a MAPF instance is considered legal if each subproblem remains solvable. Before discussing how to decompose a MAPF instance, we introduce how to check whether a decomposition of a MAPF instance is legal, i.e., every subproblem should have a solution that avoids conflicts with other subproblems’ paths.
We can simplify this by having each subproblem’s agents move until all previous subproblems have reached their targets, while avoiding the starts of subproblems that haven’t been solved yet. In this simplified scenario, every subproblem only needs to avoid previous subproblems’ targets and later subproblems’ starts. If we can find a solution for a subproblem under this simplified scenario, it is solvable. If we can find such a solution for every subproblem of a decomposition under the simplified scenario, the decomposition does not sacrifice the completeness of MAPF.
We check the solvability of a decomposition by verifying whether every agent of every subproblem can find a path under the simplified scenario. Essentially, the simplified scenario limits each subproblem’s agents to move only within a specific layer of the time-spatial space, where each layer has a unique timestep range. Therefore, we call our method LayeredMAPF. However, it is crucial to note that the solution under the simplified scenario may not be the only conflict-free solution, and the simplified scenario may generate solutions with extra wait actions. Thus, we will set the previous subproblem’s solution as constraints to avoid if a MAPF method supports it (serial MAPF methods). For MAPF methods that do not support setting external paths as constraints (parallel MAPF methods), we will set the previous subproblems’ targets and the later subproblems’ starts as occupied to ensure completeness.

### v2

@@ 2.4.2 Lazy constraints addition for MAPF (LaCAM)
It is noteworthy that when LaCAM encounters a known configuration, it reinserts the corresponding high-level node into the open set (which stores nodes for high-level search). This action prevents repeated appearance of configurations, thereby preventing LaCAM from avoiding conflicts with external paths (external path: path not belonging to current MAPF instance), as all agents may need to wait at times to avoid conflicts with external paths.
@@
Okumura et al. proposed LaCAM2 [18], which introduces two enhancements to LaCAM. The first enhancement is its anytime version, called LaCAM*, which converges to optimality eventually, provided that solution costs accumulate transition costs. The second enhancement adds a swap action to the successor generator of LaCAM, enabling the quick generation of initial solutions.
2.5 Independence detection
The idea of splitting a MAPF problem into multiple smaller subproblems has been explored by researchers in recent years. Standley et al.[19, 20] proposed that if the optimal paths of two agents have no conflicts, they can be solved independently. Standley[20] introduced an independence detection (ID) algorithm to decompose a group of agents into the smallest possible groups. ID begins by assigning each agent to its own group and finds an initial path for each group independently. It then attempts to find alternative paths to avoid conflicts, via classic MAPF methods. If attempts to find conflict-free paths fail, ID merges the conflicting groups. The process continues until there are no conflicts between agents from different groups.
Sharon et al.[21] proposed a continuum between CBS and ID called Meta-Agent CBS (MA-CBS). MA-CBS introduces a predefined parameter B, where conflicting agents are merged into a meta-agent and treated as a joint composite agent if the number of conflicts exceeds B. The original CBS algorithm corresponds to the extreme case where B = $\infty$ (never merge agents), while the ID framework[19, 20] represents the other extreme, where B = 0 (always merge agents when conflicts occur).
Compared to ID, our method considers the order of solving subproblems, thus our method have more possibility to decomposeMAPF instance into smaller subproblems. And our decomposition is decoupled from MAPF methods, while ID may need run multiple times of MAPF methods for an agent to find conflict-free paths. So in terms of time cost, our method has significant advantages. Since ID only considers making a group avoid conflict with one group at one time, a group may have conflict with a group it avoided before. For example, group A avoid conflicts wth group B, but A may have conflicts with B after A tries to avoid conflicts with group C. So it may be trapped into infinite loops. This phenomenon is not frequent when there are a few agents but more likely to happen if there are dense agents. In results, we compare our method with ID. And our method demonstrates significant advantages comparing to ID, in terms of max subproblem size, time cost and success rate to find conflict-free solution.
2.6 SEQ(Sequence) and DSP(Delayed Shortest Path)
Online Multi-Agent Path Finding (MAPF)[22] is the online version of MAPF where new agents appear over time. One kind of Online MAPF assumes when an agent reaches its goal it stays there, this results in a setting similar to Lifelong MAPF[23]. Another kind of Online MAPF assumes an agent disappears when reaching its goal.
For the second kind of online MAPF, non-conflicting paths can be generated by considering the agents sequentially according to a given order, taking their shortest paths from start to goal location and adding delays at the beginning of the paths to avoid conflicts. The algorithm SEQ (SEQUENCE)[24] does that but, for each agent, naively chooses a relative delay that is equal to the length of the shortest path of the agent that precedes it in the order. Hence, in SEQ, one agent at a time follows its shortest path to destination while the other agents wait. Delayed Shortest Path (DSP)[25] improve SEQ by calculating safe delays, i.e. relative delays that are possibly shorter than the shortest path of the previous agent in the order but are long enough to safely avoid conflicts.
DSP and SEQ use a given priority order for the agents. There are several types of priority order:
  • SH (Shortest path first) gives a higher priority to agents with shorter paths.
  • LH (Longer path first) is the opposite of SH and gives a higher priority to agents with longer paths. RND prioritizes the agents randomly.
  • LD (Lowest delays first) is a greedy method that prioritizes agents according to the lowest safe delay.
The similarity between our method and SEQ and DSP is that we also decompose a MAPF instance to subproblems with a priority order and introduce delays to get conflict-free solution. However, our method focuses on classic MAPF (i.e., offline MAPF), each agent stays at its start position before move and stay at its target position after finish its path. The biggest difference between offline MAPF and online MAPF is, agents staying in start or target positions may block other agents’ path. So essentially, our method focus on different problem comparing to SEQ and DSP.
3 Preliminaries
3.1 Basic definitions
@@
In this section, we present fundamental definitions of multi-agent pathfinding. Since our method is dimension-independent, the provided definitions apply to any-dimensional cell space $\mathcal{C}_{\mathcal{N}}$.
$\bullet$ Grid space
@@
Let $\mathcal{C}_{\mathcal{N}}$ ($\mathcal{N}\geq 2$) denote a finite $\mathcal{N}$-dimensional integer Euclidean space, where the size of the space is defined as $\mathcal{D}$, with $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{i},\ldots,d_{\mathcal{N}}\}$ and $d_{i}\in\mathbb{N}$. The coordinates of an element g in this space are defined as a vector $(x_{1},x_{2},\ldots,x_{i},\ldots,x_{\mathcal{N}})$, where $x_{i}\in([0,d_{i})\cap\mathbb{N})$.
$\bullet$ Cell states
@@
There are only two possible states for a cell/element in $\mathcal{C}_{\mathcal{N}}$: passable or unpassable. The set of all passable cells in $\mathcal{C}_{\mathcal{N}}$ is denoted as $\mathcal{F}$, while the set of all unpassable cells is denoted as $\mathcal{O}$. Therefore, $\mathcal{F}\cup\mathcal{O}=\mathcal{C}_{\mathcal{N}}$, $\mathcal{F}\cap\mathcal{O}=\emptyset$.
$\bullet$ Agents
@@
Assuming there are $k$ agents $A=\{a_{1},a_{2},\ldots,a_{k}\}$ in the grid space $\mathcal{C}_{\mathcal{N}}$, where each agent always occupies a passable cell. For convenience, we denote starts or target state of an agent $a_{i}$ as $S[a_{i}]$ or $T[a_{i}]$.
@@
Each agent has a unique starting cell and a targeting cell. Specifically, for all $i\in\{1,2,\ldots,k\}$ and $j\in\{1,2,\ldots,k\}$ with $i\neq j$, we have $S[a_{i}]\neq S[a_{j}]$, $T[a_{i}]\neq T[a_{j}]$, $S[a_{i}]\neq T[a_{j}]$, and $T[a_{i}]\neq S[a_{j}]$.
$\bullet$ Timestep
@@
Time is discretized into timesteps. At each timestep, every agent can either move to an adjacent cell or wait at its current cell.
$\bullet$ Path
@@
A path $p_{i}$ for agent $a_{i}$ is a sequence of cells that every pair of subsequent cells is adjacent or identical (indicating a wait action), starting at the start cell $s_{i}$ and ending at the target cell $t_{i}$. The cost of a path is the number of time steps it takes, i.e., the size of the path sequence $|p_{i}|$. We assume that the agents remain at their targets forever after completing their paths.
$\bullet$ Conflict
@@
We say that two agents have a conflict if they are at the same cell or exchange their positions at the same timestep (e.g., $p_{i}[t]==p_{j}[t+1]$ and $p_{i}[t+1]==p_{j}[t]$, where $p_{i}$ and $p_{j}$ means path of agent $a_{i}$ and $a_{j}$, $p_{i}[t]$ means $a_{i}$’s location at timestep $t$).
$\bullet$ Solution
@@
A quick method to determine whether a MAPF instance is solvable is whether every agent can find a collision-free path (i.e., pass no unpassable cells) from the grid space. It is noteworthy that passing through other agents’ starting or target cells is acceptable.
Here, we define a search path function:
Definition 1.
$search\_path(a_{i},\mathcal{C}_{\mathcal{N}},avoid\_node\_set)$: a complete method to search a path in $\mathcal{C}_{\mathcal{N}}$ that connects the start state and target state of agent $a_{i}$. The parameter $avoid\_node\_set$ refers to the set of nodes that cannot be part of the path. If a solution exists, we denote this as $search\_path(a_{i},\mathcal{C}_{\mathcal{N}},avoid\_node\_set)\neq\emptyset$.
$search\_path$ could be implemented using Best-First Search or Breadth-First Search. If a LA-MAPF instance is solvable, then for all agents $a_{i}\in A$, it holds that $search\_path(a_{i},\mathcal{C}_{\mathcal{N}},\emptyset)\neq\emptyset$.
Here is an example about how to check a MAPF instance is solvable, as shown in Fig. 2. We notice if a MAPF instance failed to meet this condition, it must be unsolvable; however, if a instance meet this condition, it might be unsolvable, an example is shown in Fig. 3. So our decomposition minimizing every subproblem’s possibility of unsolvable by making every subproblem meet this condition.
A
B
C
D
Figure 2: These figures show a solvable MAPF instance (Fig. A) and an unsolvable MAPF instance (Fig. C). $S_{1},S_{2},S_{3}$ and $T_{1},T_{2},T_{3}$ represent the start and target cells of the three agents in the instance, as follows. The paths of each agent in the instance are shown in Figures B and D, respectively. The instance in Fig. A passes the solvability check, as all its agents have a path to their target. However, the instance in Fig. C does not pass the solvability check and is unsolvable because only the 3rd agent has a path to its target, while the 1st and 2nd agents have no path to their target. Grey cells represent unpassable cells, while white cells represent passable cells, as follows.
Figure 3: This figure shows a MAPF instance that passes the mentioned solvability check but is unsolvable. $a_{1}$’s start $S_{1}$ have path to its target $T_{1}$ and $a_{2}$’s start $S_{2}$ have path to its target $T_{2}$, if another agent’s start and target are passable. However, due to $a_{2}$ always block the path from $S_{1}$ to $T_{1}$, this instance is unsolvable.
3.2 Decomposition of MAPF instance
Given a MAPF instance’s graph have $N$ nodes and $k$ agents, and the maximum length of path are $T$, the total number of solutions (including solution that have conflicts) is $N^{Tk}$, so the state space grows exponentially with agent size increases. In the worst cases, a complete MAPF algorithm need to travelsal all solutions to get a conflict-free solution. So the time cost of solve MAPF instance grows nearly exponentially as the number of agents increases[3].
@@ 3.2 Decomposition of MAPF instance
To reduce the cost of solving MAPF instance, we proposed a new technique: splitting a MAPF instance into multiple subproblems, each with fewer agents but under the same map, and solving these subproblems in a specific order. Here, we define the number of agents in each subproblem as the size of the subproblem. When solving a subproblem, it is crucial to ensure that the current paths do not conflict with the paths from previous subproblems. If every subproblem is solvable, and every subproblem’s solution doesn’t pass previous subproblem’s target and later subproblem’s start, we can combining all subproblems’ results (by adding wait action at the begining of each subproblem’s solution, related proof and examples can be found in Section 4) yields the solution of the original MAPF instance. If we can’t split the instance into subproblems that pass the mentioned solvability check, we don’t split it.
@@
As the time cost of solving a MAPF instance grows nearly exponentially with the number of agents, we need to minimize the size of each subproblem to reduce the total time cost of solving all subproblems. So the criteria for judging whether one decomposition is better than another involves sorting the subproblems’ sizes in decreasing order and comparing the sizes of the subproblems from the largest to the smallest. We define a decomposition that not worse than any other decomposition of this instance as a optimal decomposition. There might more than one optimal decomposition for a MAPF instace, when two decompositions have the same size of subproblems but subproblem contain different agents. The first smaller subproblem encountered indicates the better decomposition.
@@
For example, suppose there are two decompositions of a MAPF instance with 100 agents. The sizes of the subproblems for the first decomposition are 40, 20, 15, 14, and 11, while for the second decomposition, they are 40, 20, 19, 13, and 8. The first decomposition is considered better than the second one because the third subproblem’s size (15) in the first decomposition is smaller than the second decomposition’s third subproblem size (19).
@@
Considering all agent plans together can get better solution via coordination all agent’s path while solving them separately can’t coordinate agent of different subproblems, solving all agents together tends to finding a shorter solution compared to solving them separately. Therefore, MAPF methods with decomposition of MAPF instances tend to produce worse solutions than raw methods. However, we provide no theoretical analysis regarding the extent of this degradation in this manuscript; hence, we conduct empirical comparisons through extensive testing in Section 6.
Here we define the decomposition of a MAPF instance:
Definition 2.
Decomposition of a MAPF instance: Splitting a set of $k$ agents, $A=\{a_{1},a_{2},\dots,a_{k}\}$, where $k\geq 1$, into $m$ subsets of agents, $A_{1},A_{2},\dots,A_{m}$, where $m\geq 1$ and $A_{1}\cup A_{2}\cup\dots\cup A_{m}=A$. Each subset (i.e., subproblem) should be solvable independently without updating the solutions of the other subsets. A decomposition of a MAPF instance is considered legal if each subproblem passes the solvability check.
Before discussing how to decompose a MAPF instance, we first introduce the criteria for determining whether a decomposition is legal.
We simplify this by allowing the agents in each subproblem to start moving only after all agents in the previous subproblems have reached their target states, while avoiding the start states of subproblems that have not yet been solved. In other words, in this simplified scenario, each subproblem treats the agents from other subproblems as static obstacles (remaining at their start state if solved later than the current subproblem, or staying at their target state if solved earlier). An example is shown in Fig. 4.
If all subproblems of a decomposition pass the mentioned solvability check under this simplified scenario, we say that we have found a legal decomposition of the MAPF instance.
Figure 4: This figure illustrates three subproblems ($1,\{a_{1}\};2,\{a_{2}\};3,\{a_{3}\}$) of the MAPF instance shown in Fig. 2 A and checks whether these three subproblems pass the solvability check under the simplified scenario. As shown in the figure, the agent in each subproblem has a solution that avoids the target state of the previous subproblem and the start state of the next subproblem. Therefore, since every subproblem passes the solvability check, the instance can be decomposed into these three subproblems.
Figure 5: This figure shows the total solution generated by adding wait actions at the start for the MAPF instance in Fig. 4. In the total solution, only one agent from each subproblem (denoted as $a_{i}$) is moving (marked with a dashed-line box), while the agents from the other subproblems stay at their start or target states. Specifically, $a_{1}$ (from subproblem 1) moves when $0\leq t\leq 4$; $a_{2}$ (from subproblem 2) moves when $5\leq t\leq 9$; and $a_{3}$ (from subproblem 3) moves when $10\leq t\leq 13$.
If every subproblem passes the solvability check under this simplified scenario, and every subproblem is solved, we can obtain a conflict-free total solution by combining the solutions of each subproblem. Specifically, making an agent from a subproblem wait at the start state until the agents from previous subproblems have reached their target states in the spatial-time map ensures a conflict-free total solution. An example is shown in Fig. 5.
Figure 6: This figure shows an instance where our decomposition may cause a loss of solvability. Our decomposition may result in the following three subproblems: 1, {$a_{1}$}; 2, {$a_{2}$, $a_{3}$}; 3, {$a_{4}$}. However, the second subproblem is unsolvable in the simplified scenario, as $T_{1}$ and $S_{4}$ are occupied.
However, it is noteworthy that the mentioned solvability check is not always accurate; it may misjudge unsolvable subproblems as solvable, resulting in a loss of solvability. An example is shown in Fig. 6. Based on empirical observations, this phenomenon becomes more frequent in instances with dense agents. However, the decomposition of instances also becomes less effective in cases with many dense agents (which means the size of the largest subproblem will be close to the total number of agents, or the instance may not be split). In such cases, the side effect of loss of solvability is sometimes offset.
3.2.1 Cluster and level
Definition 3.
Here, we define two types of subproblems: cluster, a type of subproblem with no restrictions on the order of solving; and level, a type of subproblem that must be solved in a specific order.
For convenience, we denote all start or target states of a cluster $c$ and a level $l$ as $S[c]$ or $T[c]$, and $S[l]$ or $T[l]$, respectively.
A brief example of clusters and levels is shown in Fig. 7. A MAPF instance may be decomposed into multiple clusters, $C=\{c_{1},c_{2},\dots,c_{m}\}$, or multiple levels, $L=\{l_{1},l_{2},\dots,l_{n}\}$, where $1\leq m\leq k$ and $1\leq n\leq k$, with $k$ representing the total number of agents in the current instance.
To ensure that the decomposition of a MAPF instance into clusters or levels is legal, both the clusters and levels must satisfy certain conditions.
Theorem 1.
Legal condition for decomposition into clusters:
If a MAPF instance is decomposed into multiple clusters $C=\{c_{1},c_{2},\dots,c_{m}\}$, then $\forall c_{i}\in C$, the $avoid\_node\_set$ is defined as
$avoid\_node\_set=\{s\mid s\in\{\bigcup_{j=0}^{m}S[c_{j}]\cup\bigcup_{j=0}^{m}T[c_{j}]\},j\neq i\}.$
If $\forall a\in c_{i},search\_path(a,\mathcal{C}_{\mathcal{N}},avoid\_node\_set)\neq\emptyset$, then the decomposition of the MAPF instance into clusters is considered legal.
In other words, if the agents in each cluster have paths that do not pass through the start or target states of agents in other clusters, this subproblem passes the mentioned solvability check, and the decomposition of the MAPF instance into clusters is considered legal.
Proof 1.
If the agents in each cluster have paths that do not pass through the start or target states of agents in other clusters, then the decomposition is legal under the mentioned simplified scenario, and the decomposition of the MAPF instance into clusters is considered legal.
A: a MAPF instance about two clusters
B: a MAPF instance about two levels
Figure 7: Figure A and B show an example about two kinds of subproblem (cluster and level) in MAPF that can be solved independently.
For MAPF instance in Figure A, we can solve $a_{1}$ and $a_{2}$ separately and have no limitation in order of solve. If we solve $a_{1}$ first, a conflict-free solution might be $a_{1}$ = $\{(1,2)\rightarrow(1,1)\rightarrow(1,0)\}$, and $a_{2}$ = $\{(0,1)\rightarrow(0,1)\rightarrow(1,1)\rightarrow(2,1)\}$. If we solve $a_{2}$ first, a conflict-free solution might be $a_{1}$ = $\{(1,2)\rightarrow(1,2)\rightarrow(1,1)\rightarrow(1,0)\}$, and $a_{2}$ = $\{(0,1)\rightarrow(1,1)\rightarrow(2,1)\}$. So we say MAPF instance in Figure A can be decompose into two clusters, cluster 1 = $\{a_{1}\}$, cluster 2 = $\{a_{2}\}$.
For MAPF instance in Figure B, we can solve $a_{1}$ and $a_{2}$ separately but there is a limitation in order of solve. If we solve $a_{1}$ first, the solution of $a_{1}$ = $\{(1,2)\rightarrow(1,1)\}$, then $a_{2}$ find no conflict free solution while considering $a_{1}$ as dynamic obstacles. If we solve $a_{2}$ first, a conflict-free solution might be $a_{1}$ = $\{(1,2)\rightarrow(1,2)\rightarrow(1,1)\}$, and $a_{2}$ = $\{(0,1)\rightarrow(1,1)\rightarrow(2,1)\}$. So we say MAPF instance in Figure B can only be decompose into two levels, level 1 = $\{a_{2}\}$, level 2 = $\{a_{1}\}$. Agents in level 1 must be solved before level 2 to pass the solvability check condition and ensure the solvability of subproblems.
Theorem 2.
Legal condition for decomposition into levels:
If a MAPF instance is decomposed into multiple levels $L=\{l_{1},l_{2},\dots,l_{n}\}$, then for each $l_{i}\in L$, the $avoid\_node\_set$ is defined as
$avoid\_node\_set=\{s\mid s\in\{\bigcup_{j=0}^{i-1}T[l_{j}]\cup\bigcup_{j=i+1}^{m}S[l_{j}]\}\}.$
If $\forall a\in l_{i},search_{p}ath(a,\mathcal{C}_{\mathcal{N}},avoid_{n}ode_{s}et)\neq\emptyset$, then the decomposition of the MAPF instance into levels is considered legal.
In other words, if the agents in each level have paths that do not pass through the target states of agents in previous levels or the start states of agents in subsequent levels, then the decomposition into levels is considered legal.
Proof 2.
If the agents in each level have paths that do not pass through the target states of agents in previous levels or the start states of agents in subsequent levels, then the decomposition is legal under the mentioned simplified scenario, and the decomposition of the MAPF instance into levels is considered legal.
It is noteworthy that the legal requirements for decomposing into clusters and levels do not guarantee the solvability of subproblems. The legal requirements are used to avoid unsolvable subproblems that cannot pass the mentioned solvability check.
3.2.2 Connectivity graph
As our main focus is on whether an agent’s path passes through the start or target states of other agents, and to reduce the time cost of searching for paths, we divide the passable cells into two types: one being the agent’s start or target state, and the other being the remaining cells, which we refer to as free cells.
We then use connected component detection to split all free cells into subsets (which we call free cell groups), where every cell in a group has a path to another cell in the same group without passing through any agent’s start or target state, or cells from other groups.